Find the extreme values (if any) of the function f(x,y,z) = x^2 + 2y^2 subject to the constraint x^2 + y^2 -z^2 = 1.
Answer #1

Here basically we use the Langrange multiplier method.

We have function Z= f(nry) = n°+2y and the constraint +4-2=1 - z = n²+ y²1 ~ z-n?-y?+1=0. Now we use will method: the langranbn zon(2+1)=0 It n70, then d=-2. then Ly= -y (6+1)=0 yao (:: 1#-4) then L2=1+177=42 o=u t non le lx==(2+)-0. 172, then no thIt 47-4, M-0, Z=14, then in - nt y - 2²=1 ► 0+ ye ) = y=1+1 16 ☺ & Jo 4 thus the point (01 + 17, 14). Now to check maximumeThus the maximum value of f is 17/2 and the minimum value is 5/4